The frequency of a tuning fork is 350 Hz.find how many vibrations it executes while the sound produced by it travels a distance of 70 m
 

Grade:10

1 Answers

Eshan
5 years ago
Dear student,

The speed of sound is 340m/s. Hence the time taken by sound to travel 70m is\dfrac{70}{340}s
Hence number of vibrations in the tiven time period=\dfrac{t}{T}=\dfrac{70/340}{1/350}=72